# The Fraction of Clumpy Galaxies in JWST Surveys over $2<z<12$

### v1

@@ 10 Conclusions
We use JWST/NIRCam imaging from the JADES survey (Eisenstein et al., 2023a, b; Rieke et al., 2023b; Bunker et al., 2024; D’Eugenio et al., 2024) to identify clumpy star-forming galaxies over $2<z<12$. The evolution of the fraction of clumpy galaxies, $f_{\rm{clumpy}}$, with redshift and correlations with other physical properties were examined. A sample of 9,121 galaxies with stellar masses $\log\left(M_{\star}/M_{\odot}\right)\geq 8$ and $\log\left(\rm{sSFR/yr}^{-1}\right)\geq-9.5$ was selected by fitting their combined HST and JWST SEDs. Clumps were selected in a way similar to earlier studies based on HST (e.g., Guo et al., 2015, 2018; Shibuya et al., 2016; Sattari et al., 2023). The algorithm in Sattari et al. (2023) was used to find off-center clumps in rest-frame NUV images of galaxies. The completeness of the clump detection procedure was determined by injecting fake clumps into images of real galaxies. Clumps were selected to be $\geq 50\%$ complete as a function of redshift and stellar mass.
@@
    We find that $f_{\rm{clumpy}}$ increases with decreasing redshift from $\sim 20\%$ at $z>6.5$ to $\sim 80\%$ at $z\sim 2.75$ (Section 5 and Figure 7). This is similar to the relative increase in the cosmic volume-averaged star-formation density over the same range in redshift. Compared with the bulk of pre-JWST studies, $f_{\rm{clumpy}}$ from this work is higher at fixed redshift and increases more steeply with decreasing redshift.
@@
    The higher $f_{\rm{clumpy}}$ found in this work is due to the higher sensitivity of JWST compared with HST (Section 5.3 and Figure 8). We calculate $f_{\rm{clumpy}}$ for a sample of 1,227 galaxies that are detected in both HST and JWST images and compare measurements from both sets of images. When clumps are selected to have NUV luminosities that are $\geq 8\%$ those of their host galaxies, both measurements of $f_{\rm{clumpy}}$ are consistent with each other. When fainter clumps are selected, $f_{\rm{clumpy}}$ from JWST data are systematically higher, which is inconsistent with that from HST.
@@
    Off-center clumps make up $15-20\%$ of the rest-frame NUV luminosity of clumpy galaxies (Section 6 and Figure 9). There is no significant dependence on redshift or stellar mass of the host galaxy. At $z\sim 2.75$, our results are consistent with previous studies.
@@
    We examine correlations between $f_{\rm{clumpy}}$ and physical properties of galaxies. There is a strong correlation between stellar mass and $f_{\rm{clumpy}}$, such that more massive galaxies have higher $f_{\rm{clumpy}}$ (Section 7.1 and Figure 10). This is the opposite trend of most previous studies. We find a mild correlation between sSFR and $f_{\rm{clumpy}}$  at the lowest masses and little to none at higher masses (Section 7.2 and Figure 11). This is consistent with the results of earlier studies.
  5. 5.
    Our $f_{\rm{clumpy}}$ measurements are consistent with a physical scenario in which clumps form largely through violent disk instabilities at $z\lesssim 5.75$ and over $9<\log\left(M_{\star}/M_{\odot}\right)<11$ and through mergers at $z\gtrsim 5.75$ and over $8<\log\left(M_{\star}/M_{\odot}\right)<9$ (Section 8 and Figure 12). This is done by comparing $f_{\rm{clumpy}}$ measured from simulations with the observations. This physical scenario is bolstered when considering the observed morphologies, gas kinematics, and pair fractions of star-forming galaxies at redshifts and stellar masses that overlap with the sample examined in this work (Section 9).
We release a catalog of physical properties derived from SED fits, the empirical PSFs, and code to reproduce Figure 7 at https://github.com/alex-delavega/clumpy_galaxies_jades_2_z_12.
We thank the JADES team for designing and preparing their observations and releasing a rich public dataset. We also thank the STScI staff for enabling this science. This research has made use of the SVO Filter Profile Service “Carlos Rodrigo”, funded by MCIN/AEI/10.13039/501100011033/ through grant PID2023-146210NB-I00. All of the JWST data used in this paper can be found in MAST (Rieke et al., 2023a; Williams et al., 2023a), as can all of the HST data used in this paper (Faber, 2011).
@@ Appendix A Empirical Point Spread Functions
Empirical PSFs were measured using photutils EPSFBuilder routine, which is based on the technique developed by Anderson & King (2000). We adopt the parameters used to construct PSFs in photutils from Zhuang & Shen (2024), who characterized JWST/NIRCam PSFs in Cycle 0 calibration programs. In the present work, PSFs were created using bright (F277W mag <22<22), unsaturated, and isolated point sources that were visually inspected in each NIRCam bandpass. For the wide bandpasses, 16 point sources in GOODS-N and 17 in GOODS-S were used to estimate PSFs. For the medium bandpasses, six to ten point sources in either field were used, depending on the bandpass.
@@
Figure 13: Empirical point spread functions (PSFs) in 14 JWST/NIRCam bandpasses in JADES GOODS-N and GOODS-S are measured in this work and shown in this figure. In the top row are images of the PSFs in each bandpass in GOODS-S. Below it are the curves of enclosed energy as a function of radius for each PSF. This is repeated in the third and bottom rows for GOODS-N. The PSFs are estimated using the algorithm by Anderson & King (2000) and implemented in photutils (Bradley, 2023). The parameters that are used to construct the PSFs with photutils are adopted from Zhuang & Shen (2024). The curves of enclosed energy that are measured in this work agree to within 5% with those measured by Ji et al. (2024), who also calculated empirical PSFs in the JADES GOODS-S field. The enclosed energy curves from that work are shown as dashed lines in the second row.
Appendix B Measurements of the Fraction of Clumpy Galaxies from Previous Studies
In this Section, we describe how measurements of $f_{\rm{clumpy}}$ were extracted from the earlier studies listed in Table 2. Following Shibuya et al. (2016), $f_{\rm{clumpy}}$ from Lotz et al. (2006); Ravindranath et al. (2006); Conselice & Arnold (2009); Oesch et al. (2010); Law et al. (2012); Jiang et al. (2013); Kawamata et al. (2015) and Curtis-Lake et al. (2016) were measured using the fractions of galaxies undergoing mergers or exhibiting disturbed morphologies or multiple cores reported in these works. These come from the following tables, figures, or sections in these studies: Table 1 in Lotz et al. (2006); Table 2 in Ravindranath et al. (2006); Table 2 in Conselice & Arnold (2009); Section 4 in Oesch et al. (2010); Figure 17 in Law et al. (2012); Section 3.3 in Jiang et al. (2013); Section 3.1 in Kawamata et al. (2015); and Figure 17 in Curtis-Lake et al. (2016). $f_{\rm{clumpy}}$ from Bowler et al. (2017) was obtained from the fraction of mergers given in their Section 4.4.
Following Guo et al. (2015), $f_{\rm{clumpy}}$ from Elmegreen et al. (2007) was inferred using the abundances of clump cluster, spiral, and elliptical galaxies, given in their Figure 6. Other categories (chain galaxies, double nuclei, and tadpoles) were excluded as they generally have small axis ratios and would not have been included according to our sample selection (Section 3).
For all other studies listed in Table 2, $f_{\rm{clumpy}}$ measurements were given explicitly. These come from the following tables, figures, or sections in these works: Table 1 in Wuyts et al. (2012); Section 3.1 in Tadaki et al. (2014); Figure 10 in Guo et al. (2015); Figure 4 in Shibuya et al. (2016); Figure 8 in Ribeiro et al. (2017); Figure 16 in Huertas-Company et al. (2020); Figure 7 in Sattari et al. (2023); Figure 12 in Tan et al. (2024); and Section 4.1 in Harikane et al. (2025).

### v2

@@ IX Conclusions
We use JWST/NIRCam imaging from the JADES survey (Eisenstein et al., 2023b, a; Rieke et al., 2023a; Bunker et al., 2024; D’Eugenio et al., 2024) to identify clumpy star-forming galaxies over 2<z<9{2<z<9}. The evolution of the fraction of clumpy galaxies, $f_{\rm{clumpy}}$, with redshift and correlations with other physical properties were examined. A sample of 9,121 galaxies with stellar masses $\log\left(M_{\star}/M_{\odot}\right)\geq 8$ and $\log\left(\rm{sSFR/yr}^{-1}\right)\geq-9.5$ was selected by fitting their combined HST and JWST SEDs. Clumps were selected in a way similar to earlier studies based on HST (e.g., Guo et al., 2015, 2018; Shibuya et al., 2016; Sattari et al., 2023). The algorithm described in Section IV.1 was used to find off-center clumps in rest-frame NUV images of galaxies. The completeness of the clump detection procedure was determined by injecting fake clumps into images of real galaxies. Clumps were selected to be $\geq 50\%$ complete as a function of redshift and stellar mass.
@@
    We find that $f_{\rm{clumpy}}$ increases with decreasing redshift from $\sim 10\%$} at $z\sim 7.75$} to $\sim 70\%$} at $z\sim 2.75$ (Section V and Figure 8). This is similar to the relative increase in the cosmic volume-averaged star-formation density over the same range in redshift. Compared with the bulk of pre-JWST studies, $f_{\rm{clumpy}}$ from this work is higher at fixed redshift and increases more steeply with decreasing redshift.
@@
    The higher $f_{\rm{clumpy}}$ found in this work is mostly due to the higher sensitivity of JWST compared with HST (Section V.3 and Figure 9). We calculate $f_{\rm{clumpy}}$ for a sample of 1,227 galaxies that are detected in both HST and JWST images and compare measurements from both sets of images. We repeat this for another set of JWST images that share the resolutions of the HST images. $f_{\rm{clumpy}}$ from JWST shows little change with resolution, but increases significantly when fainter clumps are selected.
@@
    We examine the correlation between $f_{\rm{clumpy}}$ and the stellar masses of their host galaxies. There is a strong correlation between stellar mass and $f_{\rm{clumpy}}$, such that more massive galaxies have higher $f_{\rm{clumpy}}$ (Section VI and Figure 10). This is the opposite trend of most previous studies.
@@
    Our $f_{\rm{clumpy}}$ measurements are consistent with a physical scenario in which clumps form largely through violent disk instabilities at $z\lesssim 5.75$ and over $9<\log\left(M_{\star}/M_{\odot}\right)<11$ and through mergers at $z\sim 7.75$} and over $9<\log\left(M_{\star}/M_{\odot}\right)<10$} (Section VII and Figure 11). This is done by comparing $f_{\rm{clumpy}}$ measured from simulations with the observations. This physical scenario is bolstered when considering the observed morphologies, gas kinematics, and pair fractions of star-forming galaxies at redshifts and stellar masses that overlap with the sample examined in this work (Section VIII).
We release a catalog of physical properties derived from SED fits, the empirical PSFs, and code to reproduce Figure 8 on Zenodo under an open-source Creative Commons Attribution license: https://doi.org/10.5281/zenodo.19444334 (catalog doi:10.5281/zenodo.19444334).
We thank the anonymous referee for constructive comments which improved the paper.A.d.l.V. thanks Guochao Sun for enlightening discussions on formation of clumps in turbulent, high-redshift galaxies. A.d.l.V. also thanks Mark J. Achenbach for hitting the “submit” button on arXiv while A.d.l.V was driving in Hilo.We thank the JADES team for designing and preparing their observations and releasing a rich public dataset. We also thank the STScI staff for enabling this science. This research has made use of the SVO Filter Profile Service “Carlos Rodrigo”, funded by MCIN/AEI/10.13039/501100011033/ through grant PID2023-146210NB-I00.
All of the JWST data used in this paper can be found in MAST (Rieke et al., 2023b; Williams et al., 2023b), as can all of the HST data used in this paper (Faber, 2011).
@@ Appendix A Empirical Point Spread Functions
Empirical PSFs were measured using photutils EPSFBuilder routine, which is based on the technique developed by Anderson and King (2000). We adopt the parameters used to construct PSFs in photutils from Zhuang and Shen (2024), who characterized JWST/NIRCam PSFs in Cycle 0 calibration programs. In the present work, PSFs were created using bright (F277W mag <22<22), unsaturated, and isolated point sources that were visually inspected in each NIRCam bandpass. For the wide bandpasses, 16 point sources in GOODS-N and 17 in GOODS-S were used to estimate PSFs. For the medium bandpasses, six to ten point sources in either field were used, depending on the bandpass.
@@
Figure 12: Empirical point spread functions (PSFs) in 14 JWST/NIRCam bandpasses in JADES GOODS-N and GOODS-S are measured in this work and shown in this figure. In the top row are images of the PSFs in each bandpass in GOODS-S. Below it are the curves of enclosed energy as a function of radius for each PSF. This is repeated in the third and bottom rows for GOODS-N. The PSFs are estimated using the algorithm by Anderson and King (2000) and implemented in photutils (Bradley, 2023). The parameters that are used to construct the PSFs with photutils are adopted from Zhuang and Shen (2024). The curves of enclosed energy that are measured in this work agree to within 5% with those measured by Ji et al. (2024), who also calculated empirical PSFs in the JADES GOODS-S field. The enclosed energy curves from that work are shown as dashed lines in the second row.
Appendix B Treating Clumps as Point Sources in JWST Images
In this Appendix, we briefly describe how clumps are identified in JWST images when they are considered to be point sources. While clumps in JWST data typically have extended light profiles (Section IV.2), in order to compare $f_{\rm{clumpy}}$ measurements from HST and JWST images (Section V.3), clumps must be treated as point sources. This is because they were largely found to have light profiles consistent with point sources in earlier work based on HST.
To identify a complete sample of clumps under the assumption that they resemble PSFs, we perform the same routines to identify clumps and determine the completeness thereof as described in Sections IV.1 and IV.3, with only two exceptions as follows. First, aperture corrections are applied assuming the typical light profile of a clump is that of the PSF. Aperture corrections are 1.348, 1.307, 1.299, 1.336, and 1.523 in GOODS-N and 1.365, 1.328, 1.330, 1.356, and 1.485 in GOODS-S for the F090W, F115W, F150W, F200W, and F277W bandpasses, respectively. Second, to determine the completeness as a function of redshift and stellar mass, PSFs are inserted randomly into real images of galaxies, instead of Sérsic profiles convolved with PSFs. The resulting completeness curves are shown in Figure 13. It is in the same format as Figure 7. The optimal smoothing length is 10 pixels, the same as that when treating clumps as extended objects. Complete samples of clumps can be detected at fainter minimum fractional luminosities (as low as $2\%$) when modeled as PSFs. This is largely because aperture corrections are smaller.
Figure 13: Same as in Figure 7, except fake clumps are modeled as PSFs. The optimal smoothing length is 10 pixels (middle row), as was found when treating clumps as extended objects.